OpenStudy (anonymous):

solve 3√x-3-2=16 (radical ends above 3) a.x=6 b.x=9 c.x=39 d.none of the above solve √2x+15=√5(radical ends on top of 15) a.x=10 b.x=5 c.x=-5 d.None of the above rationalize the denominator of √-36 / (2-3i)+ (3+2i) I would really appreciate the help!

OpenStudy (callisto):

For the first question, add 2 to both sides, what do you get?

OpenStudy (anonymous):

\[3\sqrt{x+5}=16\] i think

OpenStudy (callisto):

Left side is correct, right side is not correct. You haven't add 2 to the right side

OpenStudy (anonymous):

so would it be 18 instead of = 16?

OpenStudy (callisto):

=18 instead of = 16 :) Here you got \(3\sqrt{x+5}=18\) Now, divide both sides by 3

OpenStudy (anonymous):

well the 18 will be 6 but I don't know how to go about dividing the radical

OpenStudy (callisto):

Wait.... Not right... It should be \(3\sqrt{x+3}=18\)

OpenStudy (callisto):

Sorry I overlooked something... Once again. Add 2 to both sides \[3\sqrt{x-3}-2=16\]\[3\sqrt{x-3}-2+2=16+2\]\[3\sqrt{x-3}=18\]Divide both sides by 3.\[\frac{3\sqrt{x-3}}{3}=\frac{18}{3}\]Simplify it. What do you get from here?

OpenStudy (anonymous):

would it be \[\sqrt{x}-1=6\]

OpenStudy (callisto):

Nope. You need not deal with the things in the square root for the moment. Just deal with the coefficient of the square root.

OpenStudy (callisto):

And you cannot break the square root in that way..

OpenStudy (anonymous):

so \[\sqrt{x-3}=6?\]

OpenStudy (callisto):

Yes. Square both sides now. What do you get?

OpenStudy (anonymous):

what do you mean by square them? I'm sorry I'm clearly really bad at math.

OpenStudy (callisto):

My bad :( Take square on both sides. \[\sqrt{x-3} ^2=6^2 \]Can you simplify it?

OpenStudy (anonymous):

so I think either \[\sqrt{x^2-9}=36\] or\[\sqrt{x-9}=36\]

OpenStudy (callisto):

Nope. Things in the square root won't change. \[\sqrt{x-3}^2=6^2\]\[x-3=36\] Since = \(\sqrt a ^2 = a\)

OpenStudy (anonymous):

so x=39 right?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

Thank you so much you helped me so much! thanks for being patient!

OpenStudy (callisto):

Welcome. question 2! \[\sqrt{2x+15}=\sqrt{5}\] Take square on both sides, what do you get?

OpenStudy (anonymous):

So \[\sqrt{2x+15}^2=\sqrt{5}^2\]

OpenStudy (callisto):

Yes, and then?!

OpenStudy (anonymous):

that removes the radicals right? so we'd get 2x+15=5

OpenStudy (callisto):

Yes. Next, subtract 15 from both sides.

OpenStudy (anonymous):

so 2x=-10 which equals x=-5 Am I correct?

OpenStudy (callisto):

I think it's correct! :)

OpenStudy (anonymous):

YYYYYYAYYYYYY! Thank you!

OpenStudy (callisto):

Welcome! Last one \[\frac{-36}{ (2-3i)+ (3+2i)}\] ^Is this the question?

OpenStudy (anonymous):

Yes that's the question

OpenStudy (callisto):

Uh-oh, or is it\[\frac{\sqrt{-36}}{ (2-3i)+ (3+2i)}\]

OpenStudy (anonymous):

sorry yes its the second one

OpenStudy (callisto):

\(\sqrt{-1} = i\) \(\sqrt{-36} = 6i\) So, we can get \[\frac{\sqrt{-36}}{ (2-3i)+ (3+2i)} = \frac{6i}{ (2-3i)+ (3+2i)} \]Got it so far?

OpenStudy (anonymous):

yes I get it so far

OpenStudy (callisto):

Hold on.. simplify the denominator first, can you?

OpenStudy (anonymous):

how would I do that??

OpenStudy (callisto):

(2-3i) + (3+2i) = (2+3) + (2i-3i) = ...?

OpenStudy (anonymous):

so 5+ -i

OpenStudy (callisto):

5+ (-i) = 5-i So, you get \[\frac{\sqrt{-36}}{ (2-3i)+ (3+2i)}\]\[ = \frac{6i}{ (2-3i)+ (3+2i)}\]\[ = \frac{6i}{5-i}\]Now multiply the fraction by a conjugate \( \frac{5+i}{5+i}\) \[ = \frac{6i}{5-i} \times \frac{5+i}{5+i}\]\[ = \frac{6i(5+i)}{(5-i)(5+i)} \]Can you simplify it?

OpenStudy (anonymous):

Oh okay I get it so would it be \[-6+30i \over 26\]

OpenStudy (callisto):

You can further simplify it :)

OpenStudy (anonymous):

how??

OpenStudy (callisto):

They have the common factor of 2.

OpenStudy (callisto):

\[2(-3+15i) \over 2\times13\]

OpenStudy (anonymous):

\[-3+15i \over 13\] right ??

OpenStudy (callisto):

Yup.

OpenStudy (anonymous):

Thank you big time! I truly appreciate your help and your time. Thank you again!